Composite scores on the ACT and the SAT are normally distributed. Jane scored 26 on the ACT and 1200 on the SAT. Numerical summaries are shown in the table below. Interpret the probabilities and determine on which test Jane performed better.

write 3 to 5 sentences to compose your answer:
Answer #1

score on the ACT and the SAT are normally distributed.

Here we need to calculate the z score to see whether which performance is good

for ACT

Z= \frac{1200-1026}{210}=0.83

for SAT

Z=\frac{26-21.1}{5.1}=0.96

since the Z score for SAT is high as compared to ACT we can conclude that Jane performed better in SAT